Reports 1-1 of 1 Clear search Modify search
PEM (Center)
takaaki.yokozawa - 13:37 Tuesday 10 February 2026 (36326) Print this report
Installation of the accelerometers at the PSL room
I installed two portable accelerometers near the Shaker 1 and Shaker 4.
During this work, I kept the LSC_LOCK guardian with observation state, and IFO kept locking(Turned on the light, work on the optical table).

ACC1 (near the shaker1, Left plot)
K1:PEM-ACC_PSL_PORTABLE_1_OUT (no DQ channel...)

ACC2 (near the shaker2, Right plot)
K1:PEM-ACC_PSL_PORTABLE_2_OUT (no DQ channel...)

If possible, I would like to restart the DAQ after changing the k1pemmcf0 model (add them to DQ channel)
at the early morning 2/11(Wed)
Images attached to this report
Comments to this report:
takaaki.yokozawa - 8:33 Thursday 12 February 2026 (36334) Print this report
I noticed they are DQ channel.
Search Help
×

Warning

×